# Linear Bandits with Partially Observable Features

### v2

@@ Appendix F Algorithm-Agnostic Lower Bound of Regret Ignoring Unobserved Features
In this section, extending our argument in Theorem 1, we show that there exists a problem instance where linear bandit algorithms relying solely on observed features can incur regret that grows linearly in $T$. We begin by formally defining such algorithms.
@@
For each $t\in[T]$, let $\pi_{t}:\mathbb{R}^{d}\times\mathbb{R}^{t-1}\to[0,1]$, be a policy that maps an observed feature vector $\mathbf{x}\in\{\mathbf{x}_{a}:a\in[K]\}$, given past reward observation $\{y_{a_{s},s}:s\in[t-1]\}$, to a probability of selection. Then the policy $\pi_{t}$ is dependent only on observed features if, for any $y_{a_{1},1},\ldots,y_{a_{t-1},t-1}$, it holds that $\mathbf{x}_{1}=\mathbf{x}_{2}$ implies $\pi_{t}(\mathbf{x}_{1}|y_{a_{1},1},\ldots,y_{a_{t-1},t-1})=\pi_{t}(\mathbf{x}_{2}|y_{a_{1},1},\ldots,y_{a_{t-1},t-1})$.
For instance, the UCB and Thompson sampling-based policies for linear bandits (with observed features), considered in Theorem 1, satisfy Definition 1, as they assign the same selection probability as long as the observed features are the same. In contrast, the policy in the MAB algorithms (that disregard observed features) may assign different selection probability although the observed features are equal and is not dependent on the observed features. In the theorem below, we particularly provide a lower bound for algorithms that employ policies that are dependent on the observed features.
@@
For any algorithm $\Pi:=(\pi_{1},\ldots,\pi_{T})$ that consists of policies $\{\pi_{t}:t\in[T]\}$ that are dependent on observed features, there exists a set of features $\{\mathbf{z}_{1},\ldots,\mathbf{z}_{K}\}$ and a parameter $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{d_{z}}$ such that the cumulative regret
$\mathrm{Reg}_{\Pi}(T,\theta_{\star},\mathbf{z}_{1},\ldots,\mathbf{z}_{K})\geq\frac{T}{6}.$
@@
We start the proof by providing a detailed account of the scenario described in the theorem. Without loss of generality, we consider the case where $K=3$. As stated in the theorem, $a_{\star}$ represents the index of the optimal action when considering the entire reward, including both observed and latent components. In contrast, $a_{o}$ denotes the index of the optimal action when considering only the observed components. We introduce an additional notation, $a^{\prime}$, which refers to an action whose observed features are identical to those of $a_{\star}$, but with a distinct latent component. Specifically, this implies that $a^{\prime}\neq a_{\star}$ and $\mathbf{z}_{a^{\prime}}\neq\mathbf{z}_{a_{\star}}$, but $\mathbf{x}_{a^{\prime}}=\mathbf{x}_{a_{\star}}$. By definition of the policy $\pi_{t}$ that depends on the observed features, $\pi_{t}(\mathbf{x}_{a_{\star}})=\pi_{t}(\mathbf{x}_{a^{\prime}})$ and the probability of selecting an optimal arm is $\pi_{t}(\mathbf{x}_{a_{\star}})\leq 1/2$.
Taking this scenario into account, the observed part of the features associated with $a_{\star}$, $a^{\prime}$, and $a_{o}$ are defined as follows:
@@
Additionally, we define the unobserved feature vectors for actions $a_{\star}$, $a^{\prime}$, and $a_{o}$ as follows:
@@
thus it follows that $\boldsymbol{\theta}_{\star}^{(o)}=[1/3d,\dots,1/3d]^{\top}\in\mathbb{R}^{d}$ and $\boldsymbol{\theta}_{\star}^{(u)}=[2/3d_{u},\dots,2/3d_{u}]^{\top}\in\mathbb{R}^{d_{u}}$. Note that it is straightforward to verify that $|\langle\mathbf{z}_{a},\boldsymbol{\theta}_{\star}\rangle|\leq 1$, thereby satisfying the assumption on the mean reward (Section 3.2). With this established, we can also observe that the expected reward for the three actions are defined as:
@@
At each round $t\in[T]$, for any policy $\pi_{t}$ satisfying Definition 1, we have $\pi_{t}(\mathbf{x}_{a_{\star}}|y_{a_{1},1},\ldots,y_{a_{t-1},t-1})=\pi_{t}(\mathbf{x}_{a^{\prime}}|y_{a_{1},1},\ldots,y_{a_{t-1},t-1})$. This implies $\mathbb{P}(a_{t}=a_{\star}=\mathbb{P}(a_{t}=a^{\prime})$ and
@@
and the probability of selecting an optimal arm cannot exceed $1/2$. Thus, the expected regret,
@@ G.1 A Hoeffding bound for Matrices
[…]
@@
[…]
$\mathbb{P}\left(\left\|\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right\|_{2}\geq x\right)\leq\mathbb{P}\left(\lambda_{\max}\left(\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\geq x\right)+\mathbb{P}\left(\lambda_{\max}\left(-\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\geq x\right).$
We bound the first term and the second term is bounded with similar arguement. For any $v>0$,
@@
where the last inequality holds since $\exp(v\sum_{\tau=1}^{t}\mathbf{D}_{\tau})$ has nonnegative eigenvalues. Taking expectation on both side gives,
@@
By Lieb’s theorem (Tropp, 2015) the mapping $\mathbf{D}\mapsto\exp(\mathbf{H}+\log\mathbf{D})$ is concave on positive symmetric matrices for any symmetric positive definite $H$. By Jensen’s inequality,
@@
By Hoeffding’s lemma,
$e^{vx}\leq\frac{b-x}{2b}e^{-vb}+\frac{x+b}{2b}e^{vb},$
for all $x\in[-b,b]$. Because the eigenvalue of $\mathbf{D}_{\tau}$ lies in $[-b,b]$, we have
@@
Now we recursively upper bound Eq. 24 as follows:
@@
$\leq\text{Tr}\mathbb{E}\left[\exp\left(v\sum_{\tau=1}^{t-2}\mathbf{D}_{\tau}+(\frac{v^{2}b^{2}}{2})\mathbf{I}_{d}+\log\mathbb{E}\left[\left.\exp(vD_{t-1})\right|\mathcal{F}_{t-2}\right]\right)\right]$
@@ G.2 A Bound for the Gram Matrix
The Hoeffding bound for matrices (Lemma 2) implies the following bound for the two Gram matrices $\mathbf{A}_{t}:=\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}^{\top}$ and $\mathbf{V}_{t}:=\sum_{\tau=1}^{t}\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}$
@@
For any $\epsilon\in(0,1)$ and $t\geq 8\epsilon^{-2}(1-p)^{-2}K^{2}\log\frac{2Kt^{2}}{\delta}$, with probability at least $1-\delta/t^{2},$
@@
$\left\|\sum_{a\in[K]}\left(\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}-1\right)\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\right\|_{2}$ $\leq\left(\frac{K-1}{1-p}+K-2\right)\max_{a\in[K]}\left\|\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\right\|_{2}$
@@
$\leq\frac{2K}{1-p}\cdot\frac{1}{t}.$ (25)
Note that the last inequality (Eq. 25) holds by Sherman-Morrison formula. By Hoeffding bound for matrix (Lemma 2), for $x>0$, it holds that:
@@
Setting $x=\epsilon\in(0,1)$ which will be determined later, for $t\geq 8\epsilon^{-2}(1-p)^{-2}K^{2}\log\frac{2Kt^{2}}{\delta}$ with probability at least $1-\delta/t^{2},$
@@ G.3 An error bound for the Lasso estimator
Let $\{\mathbf{x}_{\tau}\}_{\tau\in[t]}$ denote the covariates in $[-1,1]^{d}$ and $y_{\tau}=\mathbf{x}_{\tau}^{\top}\bar{\mathbf{w}}+e_{\tau}$ for some $\bar{\mathbf{w}}\in\mathbb{R}^{d}$ and $e_{\tau}\in\mathbb{R}$. For $\lambda>0$, let
@@
Let $\bar{\mathcal{S}}:=\{i\in[d]:\bar{\mathbf{w}}(i)\neq 0\}$ and $\boldsymbol{\Sigma}_{t}:=\sum_{\tau=1}^{t}\mathbf{x}_{\tau}\mathbf{x}_{\tau}^{\top}$. Suppose $\boldsymbol{\Sigma}_{t}$ has positive minimum eigenvalue and $\|\sum_{\tau=1}^{t}e_{\tau}\mathbf{x}_{\tau}\|_{\infty}\leq\lambda/2$. Then,
@@
The proof is similar to that of Lemma B.4 in (Kim et al., 2024), but we provide a new proof for the (unrestricted) minimum eigenvalue condition. Let $\mathbf{X}_{t}^{\top}:=(\mathbf{x}_{1},\ldots,\mathbf{x}_{t})\in[-1,1]^{d\times t}$ and $\mathbf{e}_{t}^{\top}:=(e_{1},\ldots,e_{t})\in\mathbb{R}^{t}$. We write $\mathbf{X}_{t}(j)$ and $\widehat{\mathbf{w}}_{t}(j)$ as the $j$-th column of $\mathbf{X}_{t}$ and $j$-th entry of $\widehat{\mathbf{w}}_{t}$, respectively. By definition of $\widehat{\mathbf{w}}_{t}$,
$\|\mathbf{X}_{t}\left(\bar{\mathbf{w}}-\widehat{\mathbf{w}}_{t}\right)+\mathbf{e}_{t}\|_{2}^{2}+\lambda\|\widehat{\mathbf{w}}_{t}\|_{1}\leq\|\mathbf{e}_{t}^{(j)}\|_{2}^{2}+\lambda\|\bar{\mathbf{w}}\|_{1},$
@@
$\geq$ $\sum_{i\in\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)}|-\sum_{i\in\mathcal{S}_{\star}}|{\widehat{\mathbf{w}}_{t}(i)-\bar{\mathbf{w}}(i)}|+\sum_{i\in[d]\setminus\bar{\mathcal{S}}}|{\bar{\mathbf{w}}(i)}|$
@@
where the last inequality holds by Cauchy-Schwarz inequality. Plugging in Eq. 26 gives,
$\|\mathbf{X}_{t}\left(\bar{\mathbf{w}}-\widehat{\mathbf{w}}_{t}\right)\|_{2}^{2}\leq$ $2\lambda\sum_{i\in\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)-\bar{\mathbf{w}}(i)}|$
$\leq$ $2\lambda\sqrt{\frac{|{\bar{\mathcal{S}}}|}{\lambda_{\min}(\boldsymbol{\Sigma}_{t})}}\|\mathbf{X}_{t}\left(\bar{\mathbf{w}}-\widehat{\mathbf{w}}_{t}\right)\|_{2}$

### v3

@@ Appendix F Algorithm-Agnostic Lower Bound of Regret Ignoring Unobserved Features
In this section, extending our argument in Theorem˜1, we show that there exists a problem instance such that linear bandit algorithms relying solely on observed features can incur regret that grows linearly in $T$. We begin by formally defining such algorithms.
@@
For each $t\in[T]$, let $\pi_{t}:\mathbb{R}^{d}\times\mathbb{R}^{t-1}\to[0,1]$ be a policy that maps an observed feature vector $\mathbf{x}\in\{\mathbf{x}_{a}:a\in[K]\}$, given past reward observations $\{y_{a_{s},s}:s\in[t-1]\}$, to a probability of selection. Then the policy $\pi_{t}$ is dependent only on observed features if, for any $y_{a_{1},1},\ldots,y_{a_{t-1},t-1}$, it holds that $\mathbf{x}_{1}=\mathbf{x}_{2}$ implies $\pi_{t}(\mathbf{x}_{1}|y_{a_{1},1},\ldots,y_{a_{t-1},t-1})=\pi_{t}(\mathbf{x}_{2}|y_{a_{1},1},\ldots,y_{a_{t-1},t-1})$.
For instance, the UCB and Thompson sampling-based policies for linear bandits (with observed features), considered in Theorem˜1, satisfy Definition˜1, as they assign the same selection probability as long as the observed features are the same. In contrast, policies in MAB algorithms (which disregard observed features) may assign different selection probabilities even when the observed features are equal, and thus are not dependent on the observed features. In the theorem below, we particularly provide a lower bound for algorithms that employ policies that are dependent on the observed features.
@@
For any algorithm $\Pi:=(\pi_{1},\ldots,\pi_{T})$ that consists of the policies $\{\pi_{t}:t\in[T]\}$ that are dependent on observed features, there exists a set of features $\{\mathbf{z}_{1},\ldots,\mathbf{z}_{K}\}$ and a parameter $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{d_{z}}$ such that the cumulative regret
$\mathrm{Reg}_{\Pi}(T,\boldsymbol{\theta}_{\star},\mathbf{z}_{1},\ldots,\mathbf{z}_{K})\geq\frac{T}{6}.$
@@
We start the proof by providing a detailed account of the scenario described in the theorem. Without loss of generality, we consider the case where $K=3$. As stated in the theorem, $a_{\star}$ represents the index of the optimal action when considering the entire reward, including both observed and latent components. In contrast, $a_{o}$ denotes the index of the optimal action when considering only the observed components. We introduce an additional notation, $a^{\prime}$, which refers to an action whose observed features are identical to those of $a_{\star}$, but with a distinct latent component. Specifically, this implies that $a^{\prime}\neq a_{\star}$ and $\mathbf{z}_{a^{\prime}}\neq\mathbf{z}_{a_{\star}}$, but $\mathbf{x}_{a^{\prime}}=\mathbf{x}_{a_{\star}}$. By definition of the policy $\pi_{t}$ that depends on the observed features, $\pi_{t}(\mathbf{x}_{a_{\star}})=\pi_{t}(\mathbf{x}_{a^{\prime}})$ and the probability of selecting the optimal arm is $\pi_{t}(\mathbf{x}_{a_{\star}})\leq 1/2$.
Taking this scenario into account, the observed part of the features associated with $_{\star}, aa^{\prime}$, and $a_{o}$ are defined as follows:
@@
Additionally, we define the unobserved feature vectors for actions $_{\star}, aa^{\prime}$, and $a_{o}$ as follows:
@@
thus it follows that $\boldsymbol{\theta}_{\star}^{(o)}=[1/3d,\dots,1/3d]^{\top}\in\mathbb{R}^{d}$ and $\boldsymbol{\theta}_{\star}^{(u)}=[2/3d_{u},\dots,2/3d_{u}]^{\top}\in\mathbb{R}^{d_{u}}$. Note that it is straightforward to verify that $|\langle\mathbf{z}_{a},\boldsymbol{\theta}_{\star}\rangle|\leq 1$, thereby satisfying the assumption on the mean reward (Section˜3.2). With this established, we can also observe that the expected rewards for the three actions are defined as:
@@
At each round $t\in[T]$ for any policy $\pi_{t}$ satisfying Definition˜1, we have
$\pi_{t}(\mathbf{x}_{a_{\star}}|y_{a_{1},1},\ldots,y_{a_{t-1},t-1})=\pi_{t}(\mathbf{x}_{a^{\prime}}|y_{a_{1},1},\ldots,y_{a_{t-1},t-1}),$
which implies $\mathbb{P}(a_{t}=a_{\star})=\mathbb{P}(a_{t}=a^{\prime})$ and
@@
hence the probability of selecting an optimal arm cannot exceed $1/2$. Thus, the expected regret,
@@ G.1 A Hoeffding bound for Matrices
Let $\{\mathbf{M}_{\tau}:\tau\in[t]\}$ be a $\mathbb{R}^{d\times d}$-valued stochastic process adapted to the filtration $\{\mathcal{F_{\tau}}:\tau\in[t]\}$, i.e., $\mathbf{M}_{\tau}$ is $\mathcal{F}_{\tau}$-measurable for $\tau\in[t]$. Suppose that the matrix $\mathbf{M}_{\tau}$ is symmetric and the eigenvalues of the difference $\mathbf{M}_{\tau}-\mathbb{E}[{\mathbf{M}_{\tau}}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{\tau-1}}]$ lie in [b,b][-b,b] for some $b>0$. Then for $x>0$,
@@
The proof adapts Hoeffding’s inequality to a matrix stochastic process, following the argument of Tropp (2012). Let $\mathbf{D}_{\tau}:=\mathbf{M}_{\tau}-\mathbb{E}[{\mathbf{M}_{\tau}}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{\tau-1}}]$. Then, for $x>0$,
$\mathbb{P}\left(\left\|\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right\|_{2}\geq x\right)\leq\mathbb{P}\left(\lambda_{\max}\left(\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\geq x\right)+\mathbb{P}\left(\lambda_{\max}\left(-\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\geq x\right).$ (33)
We bound the first term of Eq.˜33; a similar argument applies to the second term. For any $v>0$,
@@
where the last inequality holds since $\exp(v\sum_{\tau=1}^{t}\mathbf{D}_{\tau})$ has nonnegative eigenvalues. Taking expectations on both sides yields
@@
By Lieb’s theorem (Tropp, 2015), the map $\mathbf{D}\mapsto\text{Tr}\exp(\mathbf{H}+\log\mathbf{D})$ is concave over the set of symmetric positive definite matrices for any symmetric positive definite $\mathbf{H}$. Applying Jensen’s inequality, we obtain
@@
By convexity of $e^{vx}$ and Hoeffding’s lemma, for all $x\in[-b,b]$,
$e^{vx}\leq\frac{b-x}{2b}e^{-vb}+\frac{x+b}{2b}e^{vb}.$
Since the eigenvalues of $\mathbf{D}_{t}$ lie within [b,b][-b,b], it follows that
@@
Now we recursively upper bound Eq.˜34 as follows:
@@
$\leq\text{Tr}\mathbb{E}\left[\exp\left(v\sum_{\tau=1}^{t-2}\mathbf{D}_{\tau}+(\frac{v^{2}b^{2}}{2})\mathbf{I}_{d}+\log\mathbb{E}\left[\left.\exp(v\mathbf{D}_{t-1})\right|\mathcal{F}_{t-2}\right]\right)\right]$
@@ G.2 A Bound for the Gram Matrix
The Hoeffding bound for matrices (Lemma˜2) implies the following bound for the two Gram matrices $\mathbf{A}_{t}:=\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}^{\top}$ and $\mathbf{V}_{t}:=\sum_{\tau=1}^{t}\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}$
@@
For any $\epsilon\in(0,1)$ and $t\geq 8\epsilon^{-2}(1-p)^{-2}K^{2}\log({2Kt^{2}}/{\delta})$, with probability at least $1-\delta/t^{2},$
@@
$\left\|\sum_{a\in[K]}\left(\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}-1\right)\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\right\|_{2}$ $\leq\left(\sum_{a\in[K]}\left|\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}-1\right|\right)\max_{a\in[K]}\left\|\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\right\|_{2}$
$\leq\left(\frac{K-1}{1-p}+K-2\right)\max_{a\in[K]}\left\|\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\right\|_{2}$
@@
$\leq\frac{2K}{1-p}\cdot\frac{1}{t}.$
Note that the second inequality holds under the assumption $p\in(1/2,1)$, which implies $\phi_{a,\tau}^{-1}\leq(K-1)/(1-p)$, and the last inequality is obtained via the Sherman–Morrison formula. By Lemma˜2, for $x>0$, we have
@@
Setting $x=\epsilon\in(0,1)$, for $t\geq 8\epsilon^{-2}(1-p)^{-2}K^{2}\log({2Kt^{2}}/{\delta})$ with probability at least $1-\delta/t^{2},$
@@ G.3 An error bound for the Lasso estimator
Let $\{\mathbf{x}_{\tau}\}_{\tau\in[t]}\subset[-1,1]^{d}$ denote the covariates, and let $y_{\tau}=\mathbf{x}_{\tau}^{\top}\bar{\mathbf{w}}+e_{\tau}$ for some $\bar{\mathbf{w}}\in\mathbb{R}^{d}$ and $e_{\tau}\in\mathbb{R}$. For $\lambda>0$, consider the estimator
@@
Define $\bar{{\mathcal{S}}}:=\{i\in[d]:\bar{\mathbf{w}}(i)\neq 0\}$ and $\boldsymbol{\Sigma}_{t}:=\sum_{\tau=1}^{t}\mathbf{x}_{\tau}\mathbf{x}_{\tau}^{\top}$. Suppose that $\boldsymbol{\Sigma}_{t}$ has a strictly positive minimum eigenvalue and that $\left|\sum_{\tau=1}^{t}e_{\tau}\mathbf{x}_{\tau}\right|_{\infty}\leq\lambda/2$. Then,
@@
The proof follows a similar structure to Lemma B.4 in Kim et al. (2024), but we provide a new argument for the unrestricted minimum eigenvalue condition. Let $\mathbf{X}_{t}^{\top}:=(\mathbf{x}_{1},\ldots,\mathbf{x}_{t})\in[-1,1]^{d\times t}$ and $\mathbf{e}_{t}^{\top}:=(e_{1},\ldots,e_{t})\in\mathbb{R}^{t}$. We denote by $\mathbf{X}_{t}(j)$ the $j$-th column of $\mathbf{X}_{t}$ and by $\widehat{\mathbf{w}}_{t}(j)$ the $j$-th entry of $\widehat{\mathbf{w}}_{t}$. By definition of $\widehat{\mathbf{w}}_{t}$,
$\|\mathbf{X}_{t}\left(\bar{\mathbf{w}}-\widehat{\mathbf{w}}_{t}\right)+\mathbf{e}_{t}\|_{2}^{2}+\lambda\|\widehat{\mathbf{w}}_{t}\|_{1}\leq\|\mathbf{e}_{t}\|_{2}^{2}+\lambda\|\bar{\mathbf{w}}\|_{1},$
@@
$\geq\sum_{i\in\bar{{\mathcal{S}}}}|{\widehat{\mathbf{w}}_{t}(i)}|-\sum_{i\in\bar{{\mathcal{S}}}}|{\widehat{\mathbf{w}}_{t}(i)-\bar{\mathbf{w}}(i)}|+\sum_{i\in[d]\setminus\bar{{\mathcal{S}}}}|{\bar{\mathbf{w}}(i)}|$
@@
where the last inequality holds by Cauchy-Schwarz inequality. Plugging in Eq.˜35 gives,
$\|\mathbf{X}_{t}\left(\bar{\mathbf{w}}-\widehat{\mathbf{w}}_{t}\right)\|_{2}^{2}$ $\leq 2\lambda\sum_{i\in\bar{{\mathcal{S}}}}|{\widehat{\mathbf{w}}_{t}(i)-\bar{\mathbf{w}}(i)}|$
$\leq 2\lambda\sqrt{\frac{|{\bar{{\mathcal{S}}}}|}{\lambda_{\min}(\boldsymbol{\Sigma}_{t})}}\|\mathbf{X}_{t}\left(\bar{\mathbf{w}}-\widehat{\mathbf{w}}_{t}\right)\|_{2}$